OpenStudy (anonymous):

Solve the equation for l. P = 2l + 2w

jimthompson5910 (jim_thompson5910):

Start by subtracting 2w from both sides. Then divide both sides by 2 to completely isolate l

OpenStudy (anonymous):

The online hw wont accept the following answers p/2 -w=l l=(p-2w)/2

OpenStudy (anonymous):

already tried that and it doesnt work

jimthompson5910 (jim_thompson5910):

hmm why not try l = p/2 - w or l = -w + p/2

OpenStudy (anonymous):

ok i only have one guess left.. ugh hopefully its right

jimthompson5910 (jim_thompson5910):

the answer is either \[\Large l = \frac{P - 2w}{2}\] or \[\Large l = \frac{P}{2} - w\] of course, you can rearrange the terms